Suppose that two plates are set up in parallel configuration with the top plate with an electric potential of 200V and bottom plate with an electric potential of 77 V. The plates are separated by a distance of 5 cm.

(a) If the change in potential energy of a charge of magnitude X is 5 J when it is moved from the bottom plate to the top plate, find X.

(b) What is the direction and magnitude of the electric field between the plates?
Answer #1

Date: Page: AU = O(AV) (X) (200- 77) I 5 la 2 Coulomb X = 4x lo 200-77 (b E = (AV) = Ooos d=0.05 m E = 2460 IN/ Direction - F